If a= (2,1,-1), b=(1,-1,1) then a × b is


1
Expert's answer
2022-03-20T18:48:50-0400

The cross product of vectors is given by

"\vec a\times \vec b=\begin{vmatrix}
 \hat i & \hat j & \hat k\\
 2 & 1& -1\\
1&-1&1
\end{vmatrix}\\
= 0\hat i -3 \hat j-3\hat k=(0,-3,-3)"